Consider a system with 48-bit address that supports paging AND segmentation. The page size is 8KB and each page table entry (PTE) is 4B. A process can have up to 256 segments and each segment table entry (STE) is 8B. How large in bytes is the largest possible page table?
Answer #1

Logical address size = 48 bits => Virtual address space = \small 2^{48}\;bytes

Number of segments = 256 = \small 2^8\;bytes => Segment Number = \small log_2(2^8) = 8 bits

Size of segment = Logical address space / #segments = \small \frac{2^{48}\;bytes}{2^8} = 2^{40}\;bytes

Each segment will be divided into pages where page size = 8 KB = 8*1024 bytes = \small 2^{13}\;bytes

  • Number of pages = segment size / page size = \small \frac{2^{40}}{2^{13}} = 2^{27}
  • Page offset = \small log_2(8*1024\;bytes) = log_2(2^{13}) = 13 bits
  • Page Number = \small log_2(\#pages) = log_2(2^{27}) = 27 bits

Page table entry size = 4 bytes

Size of page table

= Number of entries in page table * page table entry size

= \small 2^{27}*4 \;bytes = 2^{29}\;bytes = \bold{512\;MB}